Value of 3^222mod11




1
Expert's answer
2022-03-19T02:37:19-0400

Since "\gcd(3,11)=1" and "\varphi(11)=10," by Euler's Theorem we get that "3^{10}=1 \mod 11."

Therefore,

"3^{222}\mod11=(3^{10})^{22}\cdot 3^2\mod 11=1^{22}\cdot 9\mod 11=9."